On a good if 25% discount is given, he will earn profit of 25% what will be profit if he gives 10% discount?
(Ans:50%)

Grade:11

2 Answers

Raju Garg
12 years ago

answere is 44.4%

sanatan sharma
12 years ago

let the marked price be x 

then discounted price is x - x/4= 3x/4 = S.P.

C.P. = (100/(100+P%)) * S.P.

C.P. = (100/125)* 3x/4

C.P. = 3x/5

if he gives 10% discount then S.P.=9x/10

then profit% = ((S.P.-C.P.)/C.P.)*100

  = [(9x/10 - 3x/5)/3x/5]*100

 =50%

 

please approve the answer